Situation 2 - Jared has decided to hold another 2 week longitudinal study and prove that people actually lost weight on Jared's Fantastic "All You Can Eat" Diet. While this study is truly unbiased, Jared decided that the statisticians he hired were actually giving him bad data, so he fired them and finished the study on his own. However, he failed to write down all the data required. Below is the data he has. Please help him in determining the missing data so he can prove without a doubt how great Jared's Fantastic "All You Can Eat" Diet will help people lose weight.

Excel:

Situation #2
Outcome
Lost Weight Did Not Lose Weight Totals
Eating Healthy Diet 22 37 Q6
Jed's All You Can Eat Diet Q7 10 Q8
918 Q9 Q10
Answers
Q6 =
Q7 =
Q8 =
Q9 =
Q10 =
Answer #1
Outcome
Lost Weight Did Not Lose Weight Total
Eating Healthy Diet 22 37 Q6=22+37=59
Jed's All You Can Eat Diet Q7=918-22=896 10 Q8=896+10=906
918 Q9=37+10=47 Q10=918+47=965
Answers
Q6= 59
Q7= 896
Q8= 906
Q9= 47
Q10= 965